OpenStudy (anonymous):

Use induction to prove that the following equality holds for all natural numbers n: the summation from r=1 to n of r(r+5)= (n(n+1)(n+8))/3

OpenStudy (anonymous):

I have proven the base case but I am having alot of trouble with the induction step.

OpenStudy (anonymous):

have you ever done a proof by induction before?

OpenStudy (anonymous):

yes, I'm just having trouble making it look like the right form during the induction step. I keep getting it incorrect

OpenStudy (anonymous):

ok my guess is the algebra is going to suck

OpenStudy (anonymous):

you want me to see if i can work it out? if i try just typing it i i will probably make a mistake

OpenStudy (anonymous):

the gist if it is going to be to show that \[(k+1)(k+6)+\frac{k(k+1)(k+8)}{3}=\frac{(k+1)(k+2)(k+9)}{3}\] did you get to that step?

OpenStudy (anonymous):

I got stuck somewhere in there haha! i got the first part but getting fromt he first part to the second I must have made a mistake

OpenStudy (anonymous):

ok well if you got there, and assuming the answer is right, the algebra cannot be that hard really, since we know what we are aiming for

OpenStudy (anonymous):

\[(k+1)(k+6)+\frac{k(k+1)(k+8)}{3}\] \[=\frac{3(k+1)(k+6)+k(k+1)(k+6)}{3}\] is a start

OpenStudy (anonymous):

also a typo,should be \[=\frac{3(k+1)(k+6)+k(k+1)(k+8)}{3}\]

OpenStudy (anonymous):

ok I will keep working then, now that i have an answer to work towards. Thanks!

OpenStudy (anonymous):

you can multiply all that mess out, but it might be easier to factor out the common factor of \(k+1\) in the numerator, because you have that factor on the right hand side as well

OpenStudy (anonymous):

you would get, working this way \[(k+1)[3(k+6)+k(k+8)]\] just up top lets see if we can make it work fast

OpenStudy (anonymous):

oh heck yes, works nicely this way \[(k+1)(3(k+6)+k(k+8)\] \[(k+1)(3k+18+k^2+8k)\] \[=(k+1)(k^2+11k+18)=(k+1)(k+2)(k+9)\] as needed not that bad after all!

OpenStudy (anonymous):

you are so awesome! thank you!

OpenStudy (anonymous):

lol it is just boring algebra, but if you did not use the common factor of \(k+1\) in the top the algebra would stink as you would have to work with cubes so i will take credit for that yw